### v2

# TSD-SR: One-Step Diffusion with Target Score Distillation for Real-World Image Super-Resolution

Abstract
Pre-trained text-to-image diffusion models are increasingly applied to real-world image super-resolution (Real-ISR) task. Given the iterative refinement nature of diffusion models, most existing approaches are computationally expensive. While methods such as SinSR and OSEDiff have emerged to condense inference steps via distillation, their performance in image restoration or details recovery is not satisfied. To address this, we propose TSD-SR, a novel distillation framework specifically designed for real-world image super-resolution, aiming to construct an efficient and effective one-step model. We first introduce the Target Score Distillation, which leverages the priors of diffusion models and real image references to achieve more realistic image restoration. Secondly, we propose a Distribution-Aware Sampling Module to make detail-oriented gradients more readily accessible, addressing the challenge of recovering fine details. Extensive experiments demonstrate that our TSD-SR has superior restoration results (most of the metrics perform the best) and the fastest inference speed (e.g. 40 times faster than SeeSR) compared to the past Real-ISR approaches based on pre-trained diffusion priors. Our code will be released at https://github.com/Microtreei/TSD-SR.
1 Introduction
Image super-resolution (ISR) [9, 8, 23, 26] aims to transform low-quality (LQ) images, which have experienced noise or blur, into clear high-quality (HQ) images. Differing from traditional ISR [6, 66], which assumes a known degradation process, real-world image super-resolution (Real-ISR) [46, 62] is designed to enhance real-world images that have suffered from complex and unknown degradations, thereby offering greater practical utility.
Generative models, particularly Generative Adversarial Networks (GANs) [11, 33, 37] and Diffusion Models (DMs) [41, 17, 39], have shown remarkable power in handling Real-ISR task. GAN-based methods leverage adversarial training, toggling between the generator and discriminator to produce realistic images. While GANs are capable of one-step inference, they are commonly hampered by issues like mode collapse and training instability [2]. Recently, Diffusion Models (DMs) have shown impressive performance in the realm of image generation [47, 21]. Their robust priors empower them to produce more realistic images with richer details than GAN-based methods [39, 41]. Some researchers [29, 60, 53, 56] have successfully leveraged pre-trained DMs for Real-ISR task. However, due to the iterative denoising nature of diffusion models [17], the Real-ISR process is computationally expensive.
Figure 1: Performance and efficiency comparison among Real-ISR methods. TSD-SR stands out for achieving high-quality restoration with the fastest speed among diffusion-based models. In contrast, existing models prioritize either speed or restoration performance. The performance of each method is benchmarked on an A100 GPU with the DRealSR dataset.
To achieve an efficient and one-step network akin to GANs, several pioneering methods that condense the iterations of diffusion models through distillation [15, 57, 12, 18] have been proposed [52, 54, 48]. Among these works, OSEDiff [52] introduced the Variational Score Distillation (VSD) loss [50] to Real-ISR task, achieving state-of-the-art (SOTA) one-step performance by leveraging prior knowledge from pre-trained models. Despite these advancements, our investigation has revealed two critical limitations associated with VSD in Real-ISR applications. (1) Unreliable gradient direction. VSD relies on a Teacher Model to provide a “true gradient direction.” However, this guidance proves unreliable in scenarios where initial ISR outputs are suboptimal. (2) Insufficient detail recovery. The VSD loss exhibits notable variation across different timesteps, and the uniform sampling strategy for $t$ poses challenges in aligning the score function with detailed texture recovery requirements. These findings highlight the need for enhanced approaches to address these issues effectively.
In this paper, we propose a novel method called TSD-SR to distill a multi-step Text-to-Image (T2I) DMs [38, 39, 10] into an effective one-step diffusion model tailored for the Real-ISR task. Specifically, TSD-SR consists of two components: Target Score Distillation (TSD) and Distribution-Aware Sampling Module (DASM). TSD incorporates our new proposed Target Score Matching (TSM) loss to compensate for the limitation of VSD loss. This significant score loss leverages HQ data to provide a reliable optimization trajectory for the distillation process. It effectively reduces the visual artifacts caused by deviant predictions from the Teacher Model. DASM is crafted to bolster detail recovery by strategically sampling low-noise samples that are distribution-based during training. This approach effectively allocates more optimization to early timesteps in a single iteration, enhancing the recovery of details.
Experiments on popular benchmarks demonstrate that TSD-SR achieves superior restoration performance (most of the metrics perform the best) and high efficiency (the fastest inference speed, 40 times faster than SeeSR) compared to the state-of-the-art Real-ISR methods based on pre-trained DMs, while requiring only a single inference step.
Our main contribution can be summarized as threefold:
  • We propose a novel method called TSD-SR to achieve one-step DMs distillation for the Real-ISR task.
  • We introduce Target Score Distillation (TSD) to provide a reliable gradient to enhance the realism of the Real-ISR model’s outputs.
  • We design a Distribution-Aware Sampling Module (DASM) specifically tailored to enhance the capability of detail restoration.
2 Related Work
GAN-based Real-ISR. Since SRGAN [26] first applied GAN to ISR, it has effectively enhanced visual quality by combining adversarial loss with perceptual loss [65, 7]. Subsequently, ESRGAN [45] introduced Residual-in-Residual Dense Block and a relativistic average discriminator, further improving detail restoration. Methods like BSRGAN [62] and Real-ESRGAN [46] simulate complex real-world degradation processes, achieving ISR under unknown degradation conditions, which enhances the model’s generalization ability. Although GAN-based methods are capable of adding more realistic details to images, they suffer from training instability and mode collapse [2].
Multi-step Diffusion-based Real-ISR. Some researches [44, 29, 53, 60, 56] in recent years have utilized the powerful image priors in pre-trained T2I diffusion models [64, 35, 39] for Real-SR tasks and achieved promising results. For example, StableSR [44] balances fidelity and perceptual quality by fine-tuning the time-aware encoder and employing controllable feature wrapping. DiffBiR [29] first processes the LR image through a reconstruction network and then uses the Stable Diffusion (SD) model [39] to supplement the details. SeeSR [53] attempts to better stimulate the generative power of the SD model by extracting the semantic information in the image as a conditional guide. PASD [56] introduces a pixel-aware cross attention module to enable the diffusion model to perceive the local structure of the image at the pixel level, while using a degradation removal module to extract degradation insensitive features to guide the diffusion process along with high-level information from the image. SUPIR [60] achieves a generative and fidelity capability using negative cues [16] as well as restoration-guided sampling, while using a larger pre-training model with a larger dataset to enhance the model capability. However, all of these methods are limited by the multi-step denoising of the diffusion model, which requires 20-50 iterations in inference, resulting in an inference time that lags far behind that of GAN-based methods.
One-step Diffusion-based Real-ISR. Recently, there has been a surge of interest within the academic community in one-step distillation techniques [34, 59, 40, 58, 31] for diffusion-based Real-ISR task. SinSR [48] leverages consistency preserving distillation to condense the inference steps of ResShift [61] into a single step, yet the generalization of ResShift and SinSR is constrained due to the absence of large-scale data training. AddSR [54] introduces the adversarial diffusion distillation (ADD) [40] to Real-ISR tasks, resulting in a comparatively effective four-step model. However, this method has a propensity to produce excessive and unnatural image details. OSEDiff [52] directly uses LQ images as the beginning of the diffusion process, and employs VSD loss [50] as a regularization technique to condense a multi-step pre-trained T2I model into a one-step Real-ISR model. However, due to the incorporation of alternating training strategies, OSEDiff may initially tend towards unreliable optimization directions, which may lead to the visual artifacts.
3 Methodology
Figure 2: Pipeline overview. We train one-step Student Model $G_{\theta}$ to map the low-quality image $x_{L}$ into a more realistic one. The noisy latent $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ sampled by DASM (Details can be found in Fig. 6.) will be fed into both the pre-trained Teacher and LoRA Model to produce the Variational Score loss. Subsequently, two Models’ predictions on $\boldsymbol{z_{t}}$ yield Target Score loss. Their weighted forms, namely TSD ( red flow), in conjunction with pixel-space reconstruction loss ( green flow), are leveraged to update the Student Model. After updating the Student Model, we employ the diffusion loss ( blue flow) to alternately update the LoRA Model.
3.1 Preliminaries
Problem Formulation. ISR problem aims to reconstruct a HQ image ${x}_{H}$ from a LQ input $x_{L}$ by training a parameterized ISR model $G_{\theta}$ on a dataset $\mathcal{D}=\{(x_{L},x_{H})_{i=1}^{N}\}$, where $N$ represents the number of image pairs. It is mathematically formulated as minimizing the following objective:
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(x_{L},x_{H})\sim\mathcal{D}}[\mathcal{L}_{Rec}(G_{\theta}(x_{L}),x_{H})$ (1)
$+\lambda\mathcal{L}_{Reg}(q_{\theta}(\hat{x}_{H}),p(x_{H}))]$
Here, $\mathcal{L}_{Rec}$ denotes the reconstruction loss, commonly measured using distance metrics like $L_{2}$ or $LPIPS$ [65]. The regularization term $\mathcal{L}_{Reg}$ enhances the realism and generalization of the ISR model’s outputs. This objective can be understood as aligning the ISR output $\hat{x}_{H}$’s distribution, $q_{\theta}(\hat{x}_{H})$, with the true distribution $p(x_{H})$ by minimizing the KL-divergence [25]:
$\min_{\theta}\mathcal{D}_{\mathrm{KL}}\left(q_{\theta}(\hat{x}_{H})\|p(x_{H})\right)$ (2)
While several studies [46, 62, 45] have employed adversarial loss to optimize this objective, they often encounter issues like mode collapse and training instability. Recent work [52] achieved the state-of-the-art results using Variational Score Distillation (VSD) as the regularization loss to minimizing this objective, which inspires our research.
Variational Score Distillation. Variational Score Distillation (VSD) [50] was initially introduced for text-to-3D generation by distilling a pre-trained text-to-image diffusion model to optimize a single 3D representation [36].
In the VSD framework, a pre-trained diffusion model, represented as $\boldsymbol{\epsilon}_{\psi}$, and its trainable (LoRA [19]) replica $\boldsymbol{\epsilon}_{\phi}$, are used to regularize the generator network $G_{\theta}$. As outlined in ProlificDreamer [50], the gradient with respect to the generator parameters $\boldsymbol{\theta}$ is formulated as follows:
$\begin{split}&\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\mathrm{VSD}}\left(\boldsymbol{\hat{z}},c_{y}\right)\\ &=\mathbb{E}_{t,\epsilon}\left[\omega(t)\left(\boldsymbol{\epsilon}_{\psi}(\boldsymbol{\hat{z}_{t}};t,c_{y})-\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y})\right)\frac{\partial\boldsymbol{\hat{z}}}{\partial\boldsymbol{\theta}}\right]\end{split}$ (3)
where $\boldsymbol{\hat{z}_{t}}=\alpha_{t}\boldsymbol{\hat{z}}+\sigma_{t}\boldsymbol{\epsilon}$ is the noisy input, $\boldsymbol{\hat{z}}$ is the latent outputted by the generator network $G_{\theta}$, $\boldsymbol{\epsilon}$ is a gaussian noise, and $\alpha_{t},\sigma_{t}$ are the noise-data scaling constants. $c_{y}$ is a text embedding corresponding to a caption that describes the input image, and $w(t)$ is a time-varying weighting function.
3.2 Overview of TSD-SR
As depicted in Fig. 2, our goal is to distill a given pre-trained T2I DMs into a fast one-step Student Model $G_{\theta}$ with Teacher Model $\boldsymbol{\epsilon}_{\psi}$ and trainable LoRA Model $\boldsymbol{\epsilon}_{\phi}$. We denote the latent outputs of the distilled model as $\boldsymbol{\hat{z}_{0}}$, and HQ latent representations as $\boldsymbol{z_{0}}$. Both $\boldsymbol{\hat{z}_{0}}$ and $\boldsymbol{z_{0}}$ will pass through our Distribution-Aware Sampling Module (DASM) to obtain distribution-based samples $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ (Sec. 3.4). We train $G_{\theta}$ by minimizing the two losses: a reconstruction loss in pixel space to compare the model outputs against the ground truth, and a regularization loss (from Target Score Distillation) to enhances the realism (Sec. 3.3). After updating the Student Model, we update the LoRA Model with diffusion loss. Finally, in Sec. 3.5, we present an overview of all the losses encountered during the training phase.
3.3 Target Score Distillation
Similar to [52], we introduce VSD loss into our work as a regularization loss to enhance the realism and generalization of the $G_{\theta}$ outputs. Upon reviewing VSD Eq. 3, $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ represents the gradient direction of the optimization on $G_{\theta}$ noisy outputs $\boldsymbol{\hat{z}_{t}}$, whereas $\boldsymbol{\epsilon}_{\psi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ corresponds to the gradient direction on noisy HQ latent representations $\boldsymbol{z_{t}}$. The overarching goal of model optimization is to harmonize the predictions from $\boldsymbol{\hat{z}_{t}}$ with those from $\boldsymbol{z_{t}}$. Nonetheless, this strategy encounters hurdles, especially in the early training phase: the quality of synthetic latents $\boldsymbol{\hat{z}_{t}}$ is not high enough for Teacher Model to give a precise prediction. As illustrated in Fig. 3, the Teacher Model finds it challenging to accurately predict the optimization direction for low-quality synthetic latents, with the cosine similarity to the ideal optimization direction only $0.2$ ($0.88$ on HQ latents). This problem can lead to severe visual artifacts, as evident in Fig. 4(a).
Figure 3: A visual comparison of the gradient direction. We set the timestep $t$ to 100 and calculated the cosine similarity between the predicted directions from the Teacher Model and the true direction (towards the HQ data). The prediction direction for $\boldsymbol{z_{t}}$ closely matches the true direction, but not for $\boldsymbol{\hat{z}_{t}}$, suggesting that suboptimal samples may lead to directional deviations.
(a) Naive
(b) MSE
(c) Ours
Figure 4: The visualization of different strategies. (a) The naive method introduces fake textures and fails to recover fine details. (b) MSE leads to over-smoothed generation result, lacking high-frequency information. (c) Our method offers the superior visual effects and fine textures.
A straightforward remedial measure is to employ a mean squared error (MSE) loss to align the synthetic latents with the ideal inputs of the Teacher Model, which are derived from HQ latents. However, as shown in Fig. 4(b), this approach has been observed to lead to over-smoothed results [13]. Our strategy, instead, is to align the predictions made by the Teacher Model on both synthetic and HQ latents, thereby encouraging greater consistency between them. The core idea is that for samples drawn from the same distribution, the real scores predicted by the Teacher Model should be close to each other. We refer to this approach as Target Score Matching (TSM):
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\mathbf{\hat{z}},\mathbf{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y}))\frac{\partial\mathbf{\hat{z}}}{\partial\theta}\right]\end{split}$ (4)
where the expectation of the gradient is computed across all diffusion timesteps $t\in\{1,\cdots,T\}$ and $\epsilon\sim\mathcal{N}(0,I)$. Equation 4 encapsulates the optimization loss for our Target Score Matching. Upon examining it in conjunction with the Eq. 3, we notice that VSD utilizes the prediction residual between a Teacher and a LoRA Model to drive gradient backpropagation. Conversely, our TSM employs the synthetic and the HQ data to produce the gradients. By blending these two strategies with hyperparameter weighting $\lambda$ and $1-\lambda$, we can construct our ultimate optimization loss, effectively unifying the strengths of both approaches to optimize the training process Eq. 5.
$\begin{split}\nabla_{\theta}\mathcal{L}_{\mathrm{TSD}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)[{\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-\\ {\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y})+\lambda({\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y})-{\epsilon}_{\phi}(\mathbf{\hat{z}}_{t};t,c_{y}))]\frac{\partial\mathbf{\hat{z}}}{\partial\theta}\Big{]}\end{split}$ (5)
where $w(t)$ is time-aware weighting function tailored for Real-ISR. Other symbols are in accordance with those previously mentioned. By introducing the prediction of pre-trained diffusion model on HQ latents, we have circumvented the issue of the model falling into the visual artifacts or over-smoothed problem, as illustrated in Fig. 4(c).
3.4 Distribution-Aware Sampling Module
In the VSD-based framework, we need to match the score functions predicted by the Teacher Model with those by the LoRA Model at different timesteps $t\sim\{0,1,\cdots,T\}$. However, for the Real-ISR problem, this matching ability at different $t$ is unbalanced, as illustrated in Fig. 5(a) . This phenomenon could be attributed to our leveraging of the low-frequency (LF) priors from LQ data ,but losing guidance from high-frequency (HF) details. The noisy samples $\hat{z}_{t}$ possesses low-frequency priors (from LQ data). These priors are easily learned by the LoRA Model, leading to more similar predictions (Fig. 5(b)) during LF restoration (Stage 1). However, In the Stage 2, the lack of HF details in $\hat{z}_{t}$ causes the LoRA Model to poorly perceive details, resulting in divergent predictions (Fig. 5(c)). Our goal is to reduce this divergence.
Figure 5: (a) The prediction errors of the VSD loss at different timesteps. The error divergence is more pronounced in early timesteps than later. This phenomenon is observed throughout the optimization process. (b) the visualization of Stage 1 prediction error. (c) the visualization of Stage 2 prediction error.
Existing method, for each iteration, only matches the score function at a single timestep sample $\hat{z}_{t}$, with uniform sampling strategy of time $t$. This leads to slow convergence or even difficulty in optimization during Stage 2, as the gradients of important timesteps are averaged out. To this end, we propose our Distribution-Aware Sampling Module (DASM). This module accumulates optimization gradients for earlier timestep samples in a single iteration, enabling the backpropagation of more gradients focused on detail optimization. As shown in Fig. 6, we first introduce noise into the synthetic latent representation $\boldsymbol{\hat{z}_{t}}=(1-\sigma_{t})\boldsymbol{\hat{z}_{0}}+\sigma_{t}\boldsymbol{\epsilon}$, where $\sigma$ is a weight factor and $\epsilon$ is a gaussian noise. Subsequently, we employ a LoRA Model to perform denoising, yielding the previous time noisy samples by Eq. 6:
$\boldsymbol{\hat{z}_{t-1}}=\boldsymbol{\hat{z}_{t}}+(\sigma_{t-1}-\sigma_{t})\cdot\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y}),$ (6)
where $\sigma_{t-1}$ and $\sigma_{t}$ are from the flow matching scheduler, and here LoRA Model has fit $\boldsymbol{\hat{z}_{0}}$’s distribution. Similarly, $\boldsymbol{z_{t-1}}$ can be obtained by denoising on the Teacher Model. Ultimately, in a single iteration, the gradients from noise samples along its trajectory can be accumulated to update the Student Model. Since the obtained samples follow the diffusion sampling trajectory and are directed towards early timestep, we effectively optimize the divergence in Stage 2.
Figure 6: Illustration of DASM. Top: The naive method that add noise to samples. Bottom: Our DASM leverages the priors of diffusion models to sample noise that more closely aligns with the true sampling trajectory. The noisy samples can all serve as inputs to the subsequent network to generate gradient backpropagation.
3.5 Training Objective
We summarize all the losses we used in our framework.
Student Model $G_{\theta}$. We train our Student Model with the reconstruction loss $\mathcal{L}_{Rec}$ and the regularization loss $\mathcal{L}_{Reg}$.
For the reconstruction loss, we use $LPIPS$ loss:
$\mathcal{L}_{Rec}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)=\mathcal{L}_{LPIPS}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right).$ (7)
For the regularization loss, we use our TSD loss Eq. 5. Therefore, the overall training objective for the Student Model $G_{\theta}$ is:
$\mathcal{L}_{Stu}=\mathcal{L}_{Rec}+\gamma\mathcal{L}_{Reg},$ (8)
where $\gamma$ is a weighting scalar.
LoRA Model $\boldsymbol{\epsilon}_{\phi}$. As stipulated by VSD, the fine-tuned replica $\epsilon_{\phi}$ must be trainable, with its training objective being:
$\mathcal{L}_{Diff}(\boldsymbol{{\hat{z}}},c_{y})=\mathbb{E}_{t,\boldsymbol{\epsilon}}[\left\|\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y})-\boldsymbol{\epsilon}^{\prime}\right\|^{2}],$ (9)
where $\boldsymbol{\epsilon}^{\prime}$ serves as the training target for the denoising network, representing Gaussian noise in the context of DDPM, and a gradient towards HQ for flow matching.
4 Experiments
4.1 Experimental Settings
Training Datasets. We adopt DIV2K [1], Flickr2K [42], LSDIR [27] and first 10K face images from FFHQ [20] for training. We use the same degradation pipeline as Real-ESRGAN [46] to synthesize LR-HR pairs.
Test Datasets. We evaluated our model on synthetic DIV2K-Val [1] dataset and two real-world datasets, RealSR [4] and DRealSR [51]. The real-world datasets comprised 128x128 LQ and 512x512 HQ image pairs. The synthetic dataset included 3,000 image pairs, which are created by downgrading the 512x512 HQ images croped from DIV2K-Val to 128x128 LQ images using the Real-ESRGAN [46] degradation pipeline.
Evaluation Metrics. For evaluating our method, we apply both full-reference and no-reference metrics. Full-reference metrics include PSNR and SSIM [49] (calculated on the Y channel in YCbCr space) for fidelity, LPIPS [65] and DISTS [7] for perceptual quality, and FID [14] for distribution comparison. No-reference metrics include NIQE [63], MANIQA [55], MUSIQ [22] and CLIPIQA [43].
Compared Methods. We categorize the test models into two groups: single-step and multi-step inference. The single-step inference diffusion models include SinSR [48], AddSR [54], and OSEDiff [52]. The multi-step inference diffusion models comprise StableSR [44], ResShift [61], PASD [56], DiffBIR [29], SeeSR [53], SUPIR [60], and AddSR [54]. Specifically, for AddSR, we have conducted comparisons between its single-step and four-step models. GAN-based Real-ISR methods [62, 5, 28, 46] are detailed in the supplementary material.
Implementation Details. All models mentioned are initialized from the Teacher Model (SD3 [10] in our work). We only train vae encoder and denoising network in the Student Model, frozen the vae decoder in order to preserve vae’s prior [24]. We utilize the default prompt for the Student Model, and prompts extracted from HQ for the Teacher and LoRA models when training. We use the AdamW optimizer [30] with a learning rate of 5e-6 for the Student Model and 1e-6 for the LoRA Model, and set rank of LoRA to 64 for both two models. The training process took roughly 48h, utilizing 8 NVIDIA V100 GPUs with a batch size of 16.
Table 1: Quantitative comparison with the state-of-the-art one-step methods across both synthetic and real-world benchmarks. The number of diffusion inference steps is indicated by ‘s’. The best results of each metric are highlighted in red.
Datasets Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ NIQE $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
OSEDiff-1s 27.35 0.7610 0.3177 0.2365 141.93 7.3053 63.56 0.5763 0.7053
AddSR-1s 27.77 0.7722 0.3196 0.2242 150.18 6.9321 60.85 0.5490 0.6188
SinSR-1s 28.38 0.7497 0.3669 0.2484 172.72 6.9606 55.03 0.4904 0.6412
DRealSR Ours-1s 26.01 0.7174 0.3092 0.2216 130.83 5.5044 66.66 0.5850 0.7345
OSEDiff-1s 23.94 0.6736 0.3172 0.2363 125.93 6.3822 67.52 0.6187 0.7001
AddSR-1s 24.79 0.7077 0.3091 0.2191 132.05 5.5440 66.18 0.6098 0.5722
SinSR-1s 26.27 0.7351 0.3217 0.2341 137.59 6.2964 60.76 0.5418 0.6163
RealSR Ours-1s 23.27 0.6874 0.2807 0.2185 113.31 5.0679 71.28 0.6346 0.7235
OSEDiff-1s 23.72 0.6109 0.3058 0.2138 26.34 5.3903 65.27 0.5838 0.6558
AddSR-1s 23.26 0.5902 0.3623 0.2123 29.68 4.7610 63.39 0.5657 0.5734
SinSR-1s 24.41 0.6018 0.3262 0.2068 35.55 5.9981 62.95 0.5430 0.6501
DIV2K-Val Ours-1s 22.12 0.5595 0.2704 0.1893 23.55 4.3730 71.38 0.6136 0.7273
4.2 Comparison with Existing Methods
Quantitative Comparisons. Tab. 1 show the quantitative comparison of our method with single-step diffusion models on three datasets. Ours achieves the best results on all metrics except PSNR and SSIM. SinSR and AddSR, as distilled versions of previous multi-step super-resolution methods, reduce inference steps but experience a corresponding decrease in performance metrics. OSEDiff introduces the VSD loss from 3D generation tasks into Real-ISR without fully addressing the substantial differences between the two domains. Therefore, the no-reference metrics for image quality are not very satisfied. In contrast, our proposed TSD-SR, tailored for Real-ISR, outperform other single-step models in terms of the vast majority of key metrics.
Tab. 2 show the quantitative comparison with multi-step models. We can draw the following conclusions: (1) TSD-SR demonstrates significant advantages over competing methods in terms of LPIPS, FID, and NIQE metrics. Additionally, it achieves performance that surpasses most multi-step models on DISTS, MUSIQ, and CLIPIQA. (2) DiffBIR, SeeSR, PASD, and AddSR exhibit better performance on the MANIQA metric, which may be attributed to the fact that multi-step models have more denoising iterations to produce rich details. (3) ResShift stands out with the highest PSNR and SSIM scores, while StableSR also shows notable performance regarding DISTS and FID metrics. However, both two models underperform in terms of no-reference metrics.
Finally, we explain the lower PSNR and SSIM metrics in our experiments. Several works [54, 60] have found that these reconstruction metrics are not well-suited for evaluating Real-ISR tasks. When a model recovers better details, it leads to lower reconstruction metrics, indicating a trade-off [3, 67, 32]. Refer to the supplementary material for detailed visual comparisons and analysis.
Table 2: Quantitative comparison with state-of-the-art multi-step methods across both synthetic and real-world benchmarks. The number of diffusion inference steps is indicated by ‘s’. The best and second best results of each metric are highlighted in red and blue, respectively.
Datasets Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ NIQE $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
StableSR-200s 28.04 0.7454 0.3279 0.2272 144.15 6.5999 58.53 0.5603 0.6250
DiffBIR-50s 25.93 0.6525 0.4518 0.2761 177.04 6.2324 65.66 0.6296 0.6860
SeeSR-50s 28.14 0.7712 0.3141 0.2297 146.95 6.4632 64.74 0.6022 0.6893
SUPIR-50s 25.09 0.6460 0.4243 0.2795 169.48 7.3918 58.79 0.5471 0.6749
PASD-20s 27.79 0.7495 0.3579 0.2524 171.03 6.7661 63.23 0.5919 0.6242
ResShift-15s 28.69 0.7874 0.3525 0.2541 176.77 7.8762 52.40 0.4756 0.5413
AddSR-4s 26.72 0.7124 0.3982 0.2711 164.12 7.6689 66.33 0.6257 0.7226
DRealSR Ours-1s 26.01 0.7174 0.3092 0.2216 130.83 5.5044 66.66 0.5850 0.7345
StableSR-200s 24.62 0.7041 0.3070 0.2156 128.54 5.7817 65.48 0.6223 0.6198
DiffBIR-50s 24.24 0.6650 0.3469 0.2300 134.56 5.4932 68.35 0.6544 0.6961
SeeSR-50s 25.21 0.7216 0.3003 0.2218 125.10 5.3978 69.69 0.6443 0.6671
SUPIR-50s 23.65 0.6620 0.3541 0.2488 130.38 6.1099 62.09 0.5780 0.6707
PASD-20s 25.68 0.7273 0.3144 0.2304 134.18 5.7616 68.33 0.6323 0.5783
ResShift-15s 26.39 0.7567 0.3158 0.2432 149.59 6.8746 60.22 0.5419 0.5496
AddSR-4s 23.33 0.6400 0.3925 0.2626 154.22 5.8959 71.49 0.6826 0.7225
RealSR Ours-1s 23.27 0.6874 0.2807 0.2185 113.31 5.0679 71.28 0.6346 0.7235
StableSR-200s 23.27 0.5722 0.3111 0.2046 24.95 4.7737 65.78 0.6164 0.6753
DiffBIR-50s 23.13 0.5717 0.3469 0.2108 33.93 4.6056 68.54 0.6360 0.7125
SeeSR-50s 23.73 0.6057 0.3198 0.1953 25.81 4.8322 68.49 0.6198 0.6899
SUPIR-50s 22.13 0.5279 0.3919 0.2312 31.40 5.6767 63.86 0.5903 0.7146
PASD-20s 24.00 0.6041 0.3779 0.2305 39.12 4.8587 67.36 0.6121 0.6327
ResShift-15s 24.71 0.6234 0.3473 0.2253 42.01 6.3615 60.63 0.5283 0.5962
AddSR-4s 22.16 0.5280 0.4053 0.2360 35.41 5.2584 70.99 0.6596 0.7593
DIV2K-Val Ours-1s 22.12 0.5595 0.2704 0.1893 23.55 4.3730 71.38 0.6136 0.7273
Qualitative Comparisons. Fig. 7 present visual comparisons of different Real-ISR methods. As shown in multi-step methods’ results, SeeSR uses degradation-aware semantic cues to leverage image generation priors, but it sometimes produces over-smooth textures. SUPIR demonstrates a notably robust generative capacity, however, the overproduction of extraneous details will lead to results that lack naturalism in image restoration (e.g. Adding excessive wrinkles at the corners of a young girl’s eyes). Under more realistic degradation conditions, PASD finds it challenging to restore the corresponding content, revealing a deficiency in the model’s robustness capabilities. Among single-step methods, SinSR introduces artifacts, possibly because its distilled pre-trained model is trained from scratch, lacking sufficient real-world priors, which results in unsatisfactory image restoration capabilities. AddSR produces over-smoothed results when using its 1-step model. OSEDiff offers improved restoration effects compared to SinSR and AddSR, yet it may fall short in terms of authenticity and naturalness when it comes to detail recovery. In contrast, our method effectively generates rich textures and realistic details with enhanced sharpness and contrast. Additional visual comparisons and results are provided in the supplementary material.
Figure 7: Visual comparisons of different Real-ISR methods. Please zoom in for a better view.
Complexity Comparisons
Table 3: Comparison of computational complexity across different DMs-based methods. The performance of each method is measured on an A100 GPU, using input images sized at 512 × 512 pixels. We have disregarded the loading time for model weights and data.
StableSR DiffBIR SeeSR SUPIR PASD ResShift AddSR AddSR SinSR OSEDiff Ours
Inference Step 200 50 50 50 20 15 4 1 1 1 1
Inference Time 12.4151 7.9637 5.8167 16.8704 4.8441 0.7546 1.0199 0.5043 0.1424 0.1650 0.1362
We assess the complexity of the state-of-the-art DM-based Real-ISR models as detailed in Tab. 3, focusing on inference time. The performance of each method is benchmarked on an A100 GPU with input images sized at 512 × 512 pixels. We have disregarded the loading time for model weights and data, and the main computation time includes: (1) text extraction time when a text extractor is used; (2) text encoder computation time if needed; (3) VAE encoding and decoding time; (4) denoising network execution time. It is evident that TSD-SR has a substantial advantage in inference time compared with multi-step models. Specifically, TSD-SR is more than 120 times faster than SUPIR, 90 times faster than StableSR, about 50 times faster than DiffBIR, more than 40 times faster than SeeSR, more than 35 times faster than PASD, and 4 times faster than ResShift. When compared with existing one-step models, our model boasts the fastest inference times. This is because we directly denoise from LQ data and use a fixed prompt.
4.3 User Study
We conduct a user study comparing our method with three other diffusion-based one-step super-resolution methods. To make the evaluation more comprehensive, we selected images from five categories—human faces, buildings, animals, vegetation and characters. A total of 50 participants were engaged in the voting process. We guided participants to evaluate the best restoration results based on the similarity to HQ image, structural similarity to the LQ image, and realism of textures and details. The results shown in the Fig. 8 indicate that our method receives a 69.2% approval rate from users. We scored 57.60% in Animals, 70.00% in Buildings, 68.80% in Human Faces, 65.20% in Vegetation, and 84.40% in Characters, surpassing other models.
Figure 8: The results of our user study. Left: Category-based user preference radar chart. Our model has won the most user favor across all categories. Right: User preference pie chart. Our approach has garnered a 69.2% user satisfaction rating.
Table 4: Ablation study of Target Score Matching loss and Distribution-Aware Sampling Module.
Datasets Method NIQE $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
DRealSR w/o TSM 5.6268 63.90 0.5749 0.6958
w/o DASM 6.0368 63.56 0.5812 0.7123
Full 5.5044 66.66 0.5850 0.7345
RealSR w/o TSM 5.2728 68.56 0.6338 0.6987
w/o DASM 5.5638 69.05 0.6273 0.7031
Full 5.0679 71.28 0.6346 0.7235
4.4 Ablation Study
To validate the effectiveness of TSM loss and DASM, we conducted ablation studies by removing them separately in our experiments. We selected no-reference metrics for comparison, as these are critical for assessing image quality. The results are presented in Tab. 4. We draw the following conclusions: (1) The absence of TSM loss and DASM affects the no-reference metrics NIQE, MUSIQ, MANIQA, and CLIPIQA. (2) The lack of TSM leads to a significant decrease in MUSIQ and CLIPIQA metrics, possibly due to the unreliable directions in VSD causing unrealistic generations. (3) The absence of DASM results in a decline in NIQE, MUSIQ and CLIPIQA, possibly due to the suboptimal optimization of details.
5 Conclusion and Limitation
We propose TSD-SR, an effective and one-step model based on diffusion prior for Real-ISR. TSD-SR utilizes the TSD to enhance the realism of images generated by the distillation model. And it leverages DASM to sample distribution-based samples and accumulate their gradients to enhance the recovery of details. Our experiments have demonstrated that TSD-SR outperforms existing one-step Real-ISR models in both performance and inference speed.
limitations. Although our model boasts excellent inference speed and restoration performance, it still has a large number of model parameters compared to the past GAN/non-diffusion approaches. In the future, we plan to employ pruning or quantization methods to compress the model parameters, striving for a lightweight and efficient Real-ISR model.
Supplementary Material
Appendix A Comparison with GAN-based Methods
We compared our method with GAN-based approaches in Tab. 5. While the GAN methods show advantages in full-reference metrics such as PSNR and SSIM, our model outperforms GAN-based methods on all no-reference metrics. Some research has found the limitations of full-reference methods in the field of image super-resolution [60, 54]. Full-reference metrics often assign low scores to visually appealing results because these results produce more details, some of which may not align with the ground truth. However, no-reference metrics assess image quality based on the individual image, without the need to forcibly align with the ground truth. Therefore, in more complex and realistic degradation scenarios, no-reference metrics may be more suitable for evaluating the results of image super-resolution. In Appendix C, we discuss the comparison between full-reference metrics and human preferences, and in the Fig. 9, we present a visual comparison with GAN-based methods. From the visualization, it can be observed that our model achieves better results in terms of texture details compared to GAN methods.
Table 5: Quantitative comparison with GAN-based methods on both synthetic and real-world benchmarks. The best results of each metric are highlighted in red.
Datasets Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ NIQE $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
BSRGAN 28.70 0.8028 0.2858 0.2143 155.61 6.5408 57.15 0.4847 0.5091
Real-ESRGAN 28.61 0.8051 0.2818 0.2088 147.66 6.7001 54.27 0.4888 0.4521
LDL 28.20 0.8124 0.2791 0.2127 155.51 7.1448 53.94 0.4894 0.4476
FeMASR 26.87 0.7569 0.3156 0.2238 157.72 5.9067 53.70 0.4413 0.5633
DRealSR Ours 26.01 0.7174 0.3092 0.2216 130.83 5.5044 66.66 0.5850 0.7345
BSRGAN 26.38 0.7651 0.2656 0.2123 141.24 5.6431 63.28 0.5425 0.5114
Real-ESRGAN 26.65 0.7603 0.2726 0.2065 136.29 5.8471 60.45 0.5507 0.4518
LDL 25.28 0.7565 0.2750 0.2119 142.74 5.9880 60.92 0.5494 0.4559
FeMASR 25.06 0.7356 0.2936 0.2285 141.01 5.7696 59.05 0.4872 0.5405
RealSR Ours 23.27 0.6874 0.2807 0.2185 113.31 5.0679 71.28 0.6346 0.7235
BSRGAN 24.58 0.6269 0.3502 0.2280 49.55 4.7501 61.68 0.4979 0.5386
Real-ESRGAN 24.02 0.6387 0.3150 0.2123 38.87 4.8271 60.38 0.5401 0.5251
LDL 23.83 0.6344 0.3256 0.2227 42.28 4.8555 60.04 0.5328 0.5180
FeMASR 22.45 0.5858 0.3370 0.2205 41.97 4.8679 57.94 0.4787 0.5769
DIV2K-Val Ours 22.12 0.5595 0.2704 0.1893 23.55 4.3730 71.38 0.6136 0.7273
Figure 9: Qualitative comparisons between TSD-SR and GAN-based Real-ISR methods. Please zoom in for a better view.
Appendix B More Visual Comparisons
In Figs. 10, 11 and 12, we provide more visual caparisons with other diffusion-based methods. Numerous examples demonstrate the robust restoration capabilities of TSD-SR and the high quality of the restored images.
Figure 10: Qualitative comparisons between TSD-SR and different diffusion-based methods. Our method can effectively restore the texture and details of the corresponding object under challenging degradation conditions. Please zoom in for a better view.
Figure 11: Qualitative comparisons between TSD-SR and different diffusion-based methods. Our method can effectively restore the texture and details of the corresponding object under challenging degradation conditions. Please zoom in for a better view.
Figure 12: Qualitative comparisons between TSD-SR and different diffusion-based methods. Our method can effectively restore the texture and details of the corresponding object under challenging degradation conditions. Please zoom in for a better view.
Appendix C Comparisons of Full-reference Metrics and Human Preference
We present additional comparative experiments to demonstrate the limitations of full-reference metrics in Figure 13. It can be observed that GAN-based methods with advantages in PSNR and SSIM often result in blurred details and textures. In contrast, our results offer better visual effects, yet they do not hold an advantage in terms of PSNR and SSIM metrics. This phenomenon has also been discussed in other research works [3, 67, 32], with the majority of studies suggesting that for models with strong generation and restoration capabilities, these full-reference metrics do not adequately reflect image quality. We anticipate the development of better full-reference metrics in the future to assess advanced Real-ISR methods.
Figure 13: Comparisons between full-reference metric assessments and human visual preference. Despite scoring lower on full-reference metrics, TSD-SR generates images that align with human preference.
Appendix D Theory of Target Score Matching
The core idea of Target Score Matching (TSM) is that for samples drawn from the same distribution, the real scores predicted by the Teacher Model should be close to each other. Thus we minimize the MSE loss between the Teacher Model’s predictions of $\mathbf{\hat{z}}_{t}$ and $\mathbf{z}_{t}$ by
$\begin{split}&\mathcal{L}_{\mathrm{MSE}}(\mathbf{\hat{z}},\mathbf{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)\|\epsilon_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-\epsilon_{\psi}(\mathbf{z}_{t};t,c_{y})\|_{2}^{2}\right]\end{split}$ (10)
where the expectation of the gradient is computed across all diffusion timesteps $t\in\{1,\cdots,T\}$ and $\epsilon\sim\mathcal{N}(0,I)$.
To understand the difficulties of this approach, consider the gradient of
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{MSE}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)\cdot\underbrace{\frac{\partial{\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})}{\partial\mathbf{\hat{z}}_{t}}}_{\text{Diffusion Jacobian}}\\ &\underbrace{({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y}))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\mathbf{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (11)
where we absorb $\frac{\partial\mathbf{\mathbf{\hat{z}}}_{t}}{\partial\mathbf{\mathbf{\hat{z}}}}$ and the other constant into $w(t)$. The computation of the Diffusion Jacobian term is computationally demanding, as it necessitates backpropagation through the Teacher Model. DreamFusion [36] found that this term struggles with small noise levels due to its training to approximate the scaled Hessian of marginal density. This work also demonstrated that omitting the Diffusion Jacobian term leads to an effective gradient for optimizing. Similar to their approach, we update Eq. 11 by omitting Diffusion Jacobian:
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\\ &\mathbb{E}_{t,\epsilon}\Big{[}w(t)\underbrace{({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};c_{y},t)-{\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\mathbf{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (12)
The effectiveness of the method can be proven by starting from the KL divergence. We can use a Sticking-the-Landing style gradient by thinking of ${\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t)$ as a control variate for $\hat{\epsilon}$ For detailed proof, refer to Appendix 4 of DreamFusion [36]. It demonstrates that the gradient of this loss yields the same updates as optimizing the training loss $\mathcal{L}_{\mathrm{MSE}}$ Eq. 10, excluding the Diffusion Jacobian term.
Compared with the VSD loss, we find that the term “Prediction Residual” has changed, and the two losses are similar in the gradient update mode. Specifically, we find that VSD employs identical inputs for both the Teacher and LoRA models to compute the gradient, while here TSM uses high-quality and suboptimal inputs for the Teacher Model. The losses are related to each other through $\epsilon_{\phi}(\mathbf{\hat{z}}_{t};t,c_{y})$.
Appendix E Algorithm
Algorithm 1 details our TSD-SR training procedure. We use classifier-free guidance (cfg) for the Teacher Model and the LoRA Model. The cfg value is set to 7.5.
Input: $\mathcal{D}$ = $\{x_{L},x_{H},c_{y}\}$ , pre-trained Teacher Diffusion Model including VAE encoder $E_{\psi}$, denoising network $\epsilon_{\psi}$ and VAE decoder $D_{\psi}$, the number of iterations $N$ and step size $s$ of DASM.
Output: Trained one-step Student Model $G_{\theta}$.
1 Initialize Student Model $G_{\theta}$, including $E_{\theta}\leftarrow E_{\psi}$ with trainable LoRA, $\epsilon_{\theta}\leftarrow\epsilon_{\psi}$ with trainable LoRA, $D_{\theta}\leftarrow D_{\psi}$.
2Initialize LoRA diffusion network $\epsilon_{\phi}\leftarrow\epsilon_{\psi}$ with trainable LoRA.
3while train do
4       Sample $(x_{L},x_{H},c_{y})\sim\mathcal{D}$
      /* Network forward */
5
6      $\boldsymbol{\hat{z}}\leftarrow\epsilon_{\theta}(E_{\theta}(x_{L}))$, $\boldsymbol{z}\leftarrow E_{\psi}(x_{H})$
7      $\hat{x}_{H}\leftarrow D_{\psi}(\boldsymbol{\hat{z}})$
      /* Compute reconstruction loss */
8
9      $\mathcal{L}_{Rec}\leftarrow LPIPS(\hat{x}_{H},x_{H})$
      /* Compute regularization loss */
10
11      Sample $\epsilon$ from $\mathcal{N}(0,I)$, $t$ from $\{50,\cdots,950\}$
12      $\sigma_{t}\leftarrow$ FlowMatchingScheduler($t$)
13      $\boldsymbol{\hat{z}}_{t}\leftarrow\sigma_{t}\epsilon+(1-\sigma_{t})\boldsymbol{\hat{z}}$, $\boldsymbol{z}_{t}\leftarrow\sigma_{t}\epsilon+(1-\sigma_{t})\boldsymbol{z}$
14      $\mathcal{L}_{Reg}\leftarrow\mathcal{L}_{TSD}(\boldsymbol{\hat{z}}_{t},\boldsymbol{z}_{t},c_{y})$ // Eq. 5
15      for $i\leftarrow 1$ to $N$ do
16             $cur\leftarrow t-i\cdot s$
17            $pre\leftarrow t-i\cdot s+s$
18            $\sigma_{cur}\leftarrow$ FlowMatchingScheduler($cur$)
19            $\sigma_{pre}\leftarrow$ FlowMatchingScheduler($pre$)
20            $\boldsymbol{\hat{z}}_{cur}\leftarrow\boldsymbol{\hat{z}}_{pre}+(\sigma_{cur}-\sigma_{pre})\cdot\epsilon_{\phi}(\boldsymbol{\hat{z}}_{pre};pre,c_{y})$ // Eq. 6
21            $\boldsymbol{z}_{cur}\leftarrow\boldsymbol{z}_{pre}+(\sigma_{cur}-\sigma_{pre})\cdot\epsilon_{\psi}(\boldsymbol{z}_{pre};pre,c_{y})$
22            $\mathcal{L}_{Reg}$ += $weight\cdot\mathcal{L}_{TSD}(\boldsymbol{\hat{z}}_{cur},\boldsymbol{z}_{cur},c_{y})$
23       end for
24
25      $\mathcal{L}_{G}\leftarrow\mathcal{L}_{Rec}+\gamma\mathcal{L}_{Reg}$
26      Update $\theta$ with $\mathcal{L}_{G}$
      /* Compute diffusion loss for LoRA Model */
27
28      Sample $\epsilon$ from $\mathcal{N}(0,I)$, $t$ from $\{50,\cdots,950\}$
29      $\sigma_{t}\leftarrow$ FlowMatchingScheduler($t$)
30      $\boldsymbol{\hat{z}}_{t}\leftarrow\sigma_{t}\epsilon+(1-\sigma_{t})stopgrad(\boldsymbol{\hat{z}})$
31      $\mathcal{L}_{Lora}\leftarrow\mathcal{L}_{Diff}(\boldsymbol{\hat{z}}_{t},c_{y})$ // Eq. 9
32      Update $\phi$ with $\mathcal{L}_{Lora}$
33 end while
Algorithm 1 TSD-SR Training Procedure